# Cite Pretrain: Retrieval-Free Knowledge Attribution for Large Language Models

### v1

@@ Abstract
Figure 1: CitePretrain Framework. We construct a diverse corpus comprising Wikipedia, ArXiv, Common Crawl, and novel documents for LLMs to index. Each document is indexed through a combination of passive indexing—appending a document identifier at the end of the text—and active indexing. Active indexing includes: (1) Forward augmentation, where entity-based question-answer pairs are generated from the same document to learn mappings from identifiers to facts; and (2) Backward augmentation, where related documents are retrieved and synthesized into multi-source instructions and answers with citations, enabling fact-to-identifier supervision. The model is then continually pre-trained on this corpus and further instruction-tuned. Finally, it is evaluated on downstream citation QA tasks, spanning both long-form and short-form settings.
@@ 1 Introduction
Large Language Models (LLMs) can improve the trustworthiness of their outputs by providing citations—references that justify their answers [1, 2, 3]. While LLMs are often capable of recalling training data [4], their attributions are typically unreliable under standard training, leading to hallucinated or misattributed content [5, 6]. To address this, most existing systems apply retrieval-augmented generation (RAG): they either query an external retriever, condition on the returned documents during generation [7, 8], or generate answers first and align them with relevant documents afterward [9, 10]. While effective, RAG depends on external infrastructure (e.g., search engines) and incurs additional test-time computation, and may introduce irrelevant or even conflicting information that interferes with the model’s internal knowledge [11, 12], leading to misinterpretation of the model’s reasoning [13, 14].
In this work, we explore whether LLMs can be trained to reliably cite the specific documents they encountered during the (continual) pretraining stage without retrieval. By revising the pretraining process, we aim to unlock reliable internal citation capability, enabling LLMs to ground their outputs to verifiable sources embedded in their weights. This creates a fully end-to-end experience without introducing test-time overhead. Additionally, it offers a novel form of explainability, revealing the direct provenance of the generated content within the model’s training history.
A recent study has demonstrated that, in the continual pretraining stage, models can learn to associate facts with document identifiers [15]. However, their experiments are limited to a synthetic biography dataset with uniformly structured facts about fictional individuals, failing to capture the complexity of real-world documents—which are longer, more diverse, and contain interdependent, variably expressed facts. Moreover, their focus on single-fact citation overlooks the need for multi-fact citation in realistic long-form generation tasks. To address this, we introduce a new benchmark, CitePretrainBench, which emphasizes both document complexity and long-form generation with citation—requiring models to produce extended answers while correctly attributing knowledge sources. In this benchmark, the model is required to index a mixture of (i) real-world pretraining corpora, such as Wikipedia, Common Crawl, and arXiv, and (ii) novel documents not memorized during pretraining, including fictional narratives. This setup allows us to evaluate the model’s ability not only to relearn and index known knowledge, but also to acquire and cite new knowledge from unseen sources. In the subsequent question answering phase, we include both short-form questions, which require a single fact and citation, and long-form questions, where the model should synthesize information from multiple documents to generate coherent, well-cited answers.
We train our model using a two-stage framework: an initial continual pretraining phase where the model learns and indexes both existing and new knowledge, followed by an instruction tuning phase designed specifically to elicit the desired citation behavior. During generation on downstream tasks, we constrain the citation decoding space to titles from the corpus, ensuring that all citations are valid and verifiable. Our experiments begin with a passive indexing baseline, where a document identifier is appended to each document. While this setup enables the model to copy identifiers for memorized quotes, it reveals two key limitations: 1. Complex facts $\neq$ quotes. Many evaluation questions require the synthesis or paraphrasing of information distributed across a document. The model rarely learns to associate such non-verbatim facts with the correct document identifier of the original text. 2. Granularity alone isn’t enough. Inserting the identifiers closer to each fact [15] (e.g., after every sentence or paragraph) yield marginal improvement in recall. The model still struggles to link paraphrased or compositional facts back to their original sources.
@@
To overcome these limitations, we propose Active Indexing—a strategy that goes beyond memorization and trains the model to actively incorporate identifiers across diverse contexts. During continual pretraining, we generate synthetic QA pairs that (1) restate each fact using varied linguistic forms (e.g., definitions, comparisons, summaries), and (2) train the model to either recall knowledge given a document ID or attribute the correct ID based on the fact. This yields two complementary training objectives: 1. Source → Fact (forward): Answer questions conditioned on the given document identifier, promoting internal memory retrieval and reasoning. 2. Fact → Source (backward): Predict the document identifier for a generated answer, reinforcing attribution and source grounding.
@@
Our results on Qwen-2.5-7B and 3B highlight several key findings: 1. Both forward and backward active indexing significantly outperform the baselines, and combining them yields even greater citation gains—improving performance by up to 30.2% over passive indexing. 2. Model size has a notable impact: the difference in citation precision between the 7B and 3B models is much larger than their difference in answer correctness, indicating that citation benefits more from scale. 3. Wikipedia-based citations are easier to predict compared to other tasks. Even without any indexing, models achieve strong citation performance, suggesting that titles are often memorized during pretraining. This may be due to the topic coherence and entity-centric nature of Wikipedia titles, which makes them easier to associate with content. In our ablation study, we identify two key reasons why Active Indexing is effective: (1) it presents facts in greater quantity and more diverse formats, and (2) it explicitly trains the model to utilize document identifiers. Moreover, Active Indexing continues to benefit from scaling—showing no signs of saturation even when the amount of augmented data reaches 16× the original corpus. It improves both the model’s ability to memorize document identifiers and to generalize their usage to downstream tasks.
@@ 2 Related Work
Attribution via Retrieval-Augmented Generation (RAG).
A common approach to knowledge attribution in LLMs is to generate citations that point to supporting evidence retrieved at inference time [1, 13, 16]. Such systems typically incorporate retrieval either before generation [7, 8, 17, 18] or afterward [9, 10]. While effective, this pipeline incurs additional computational overhead and may introduce inconsistencies, as the reasoning based on retrieved documents can differ from the original generation process without retrieval [11, 12, 19, 20, 21]. Fundamentally, RAG-based attributions are limited to the retrieved evidence alone and do not reflect the broader parametric knowledge encoded during pretraining. Our approach provides direct attribution by revealing the internal sources that contribute to generation.
Internal Knowledge and Memorization.
Several studies have explored the extent to which LLMs can memorize and recall training data when providing references [5, 6]. On eliciting this capability post-training, past works have explored prompting [22, 23], constrained decoding [24], and fine-tuning technique [25] for more accurate attribution. However, these methods do not provide structured mechanisms for attribution tied to pretraining process.
Source-Aware Training.
Most relevant to our work is the recently introduced source-aware training, which explicitly attaches document identifiers to training data to improve attribution [15]. While promising, prior efforts have so far been constrained to small-scale synthetic corpora and modest-size models (e.g., 1B parameters), limiting their generalizability to real-world settings. Moreover, their approach of appending the document identifier after every fact impairs the model’s language modeling ability, reducing generalization to downstream tasks. Additionally, prior evaluations have focused only on short-form citation tasks involving a single fact and a single reference, which do not reflect the complexity of realistic multi-fact, long-form generation.
We include more related works regarding generative retrieval and data attribution in Appendix B for interested readers.
@@ 3 Set-up
In this section, we describe the problem formulation, datasets, and key evaluation metrics we used.
@@ 3.1 Problem Formulation
We study whether LLMs can cite documents drawn from the corpus $\mathcal{D}$ on which it is continually pretrained. Let
$\mathcal{D}=\{(c_{i},t_{i})\}_{i=1}^{N},$
where $c_{i}\in\Sigma^{*}$ is the full text of document $i$, and $t_{i}\in\mathcal{T}$ is a unique human-readable title that serves as its identifier. The objective is to enable the LLM to learn and utilize an internal index of document contents and their identifiers during continual pretraining, and to generate factually accurate responses with precise citations during downstream question-answering tasks. We formalize this process into:
Stage 1: Continual pretraining & index learning
During continual pretraining the model should
1. absorb the novel facts in $\mathcal{D}$, and 2. learn an internal index that maps any factual span $s\subset c_{i}$ to its title $t_{i}$. To achieve this, Passive Indexing structures pretraining examples as $(c_{i},t_{i})$ pairs, where the model predicts $t_{i}$ given $c_{i}$.
Stage 2: Question answering with citation
At inference the model receives a question $q$ and produces a response $\mathcal{R}$ with citation(s)
$\mathcal{R}=\bigl{\langle}(s_{1},C_{1}),\dots,(s_{m},C_{m})\bigr{\rangle},$
where each $s_{k}\in\Sigma^{*}$ is a factual statement that answers part of $q$, and $C_{k}\subseteq\mathcal{T}$ is a set of titles whose documents entail $s_{k}$. To enable this citation behavior, we perform citation-instruction tuning on a supervised dataset to train the model to emit $(s_{k},C_{k})$ pairs. During inference, we constrain the citation decoding space to the known title set $\mathcal{T}$ to ensure that all generated citations are valid and verifiable. Evaluation focuses on two aspects: (i) the factual accuracy and relevance of the generated statements ${s_{k}}$, and (ii) the quality of citation alignment, measured by the precision and recall of the predicted title sets $C_{k}$ against gold supports.
@@ 3.2 Datasets
To study how language models index and cite documents during continual pre-training, we construct a diverse yet computationally feasible corpus that approximates realistic pre-training distributions. Our corpus includes documents from Wikipedia, Common Crawl, and scientific papers from arXiv, mirroring the representative sources commonly included in pre-training corpora. Additionally, we include entirely novel documents—i.e., sources the model has not encountered during pre-training—to evaluate the model’s ability to learn and index new knowledge. To evaluate citation performance, we use both long-form and short-form QA benchmarks, each grounded in different parts of our corpus:short-form QA tasks. Specifically: ASQA is a long-form factoid QA dataset with ambiguous questions requiring multi-document aggregation [26]. Ground-truth sources are drawn from the KILT knowledge base [27], based on the 2019/08/01 Wikipedia snapshot. ELI5 is a long-form QA dataset [28], featuring open-ended Reddit questions (e.g., how/why/what) from the “Explain Like I’m Five” forum. We retrieve supporting documents from the August 2019 Common Crawl snapshot, preprocessed using CCNet [29]. To ensure unique identifiers, we perform title deduplication by renaming duplicates with an LLM until all titles are distinct. SciQAG is a short-form science QA dataset built from scientific papers [30]. Each question-answer pair is grounded in a source document. Then we retrieve titles from arXiv. RepliQA is a short-form QA dataset based on fictional documents published after the training cutoff of most LLMs [31]. Each instance includes a synthetic document, associated questions, and gold answers. We also perform LLM-based title deduplication to ensure unique identifiers.
To keep computation manageable for academic research, we first construct a minimal continual pre-training corpus containing only the essential documents required to answer the questions in the above datasets. This core corpus totals approximately 392M tokens, and its statistics are shown in Table 1. We also perform title deduplication across the corpus to ensure identifier uniqueness. The instruction tuning data consists of examples sampled from the training sets of the target datasets. For details on corpus construction, dataset selection, and data splits, see Appendix A.
@@ 3.3 Metrics: Correctness
We first evaluate the generation’s informativeness and utility—that is, its correctness with respect to the question. For ASQA, We compute Exact Match Recall [26], which measures the recall of correct short answers by checking whether each reference answer appears as an exact substring in the model’s output. For ELI5, we use the Claim Recall to assess the correctness of generated answers—that is, how many gold claims are supported by the answer [28]. Specifically, we compute entailment scores over three sub-claims extracted from each gold answer, providing a more accurate measure of correctness. For SciQAG, we follow prior work [30] that uses LLMs (GPT-4.1 in our case) to rate answers on a 1–5 scale across multiple dimensions, with scores normalized. We adopt the Accuracy dimension, which measures how well the answer aligns with facts from the source paper, ensuring that all claims are supported by evidence. For RepliQA, we find the recall metric from [31] insufficiently informative and instead adopt the relaxed variant of FreshEval [32]—a lightweight auto-rater that uses few-shot prompting with an LLM(GPT-4.1 in our case) to evaluate answer correctness.
@@ 3.4 Metrics: Citation Quality
For long-form QA tasks (e.g., ASQA and ELI5), where answers contain multiple facts and lack a single gold reference, we follow [14] and use an NLI model (TrueTeacher; 33) to check if cited documents entail the generated claims. Citation precision is the proportion of citations that support their claims; recall is the proportion of claims that are fully supported. For short-form QA tasks (e.g., SciQAG and RepliQA), each answer corresponds to a single fact and a unique gold document. We compare model citations to the gold reference, defining precision as the fraction of citations that match, and recall as 1 if the gold citation appears, 0 otherwise. This may underestimate true accuracy, as some citations may entail the answer without matching the gold. See Appendix D for details.
Source #Docs #Tokens / Doc #Tokens Total Task Task Type
Wikipedia 30,025 3,683 110,594,287 ASQA Long-form
Arxiv 22,743 5,013 114,011,325 SciQAG Short-form
Common Crawl 32,124 4,852 155,890,988 Eli5 Long-form
RepliQA 7,182 1,236 8,880,434 RepliQA Short-form
Others 7,593 291 2,210,346 - -
Table 1: Statistics of the CitePretrain Corpus

### v2

@@ Abstract
Figure 1: CitePretrain Framework. We construct a diverse corpus (Wikipedia, ArXiv, Common Crawl, and novel documents) for LLMs to index. Each document is indexed via passive indexing (appending a document ID) and active indexing, which includes: (1) Forward augmentation: generating entity-based QA pairs to map IDs to facts; and (2) Backward augmentation: retrieving related documents to synthesize multi-source QA pairs with citations, mapping facts to IDs. The model is continually pre-trained and instruction-tuned, then evaluated on long- and short-form citation QA tasks.
@@ 1 Introduction
Large Language Models (LLMs) can improve the trustworthiness of their outputs by providing citations—references that justify their answers (Rashkin et al., 2023; Wang et al., 2023; Huang et al., 2024b). However, references directly generated by standalone LLMs (i.e, internal citations) are unreliable (Agrawal et al., 2024), with hallucination rates of 86% Zuccon et al. 2023 and up to 91.4% (Chelli et al., 2024), and misattribution rates of 24–46% even among the few authentic ones (Walters & Wilder, 2023; Bhattacharyya et al., 2023). To address this, most existing systems apply external citations by querying an external retriever: they either condition on the retrieved documents during generation(Nakano et al., 2021; Menick et al., 2022b; Gao et al., 2023c), or align answers with documents afterward (He et al., 2023; Gao et al., 2023b).
While effective, this approach carries both practical and explainability limitations. On the practical side, it adds inference overhead from long contexts (Liu et al., 2024) and extra query optimization (Song & Zheng, 2024), depends on external infrastructure (e.g., web search) whose results can be volatile (Fang et al., 2025), and can degrade reasoning fidelity when retrieved context misses or conflicts with parametric knowledge (Xie et al., 2024; Huang et al., 2025). Moreover, many questions are answerable directly from parametric memory (Mallen et al., 2023), making these costs unnecessary. On the explainability side, external retrieval offers limited insight into what the model internally knows or recalls. Internal citation offers a potential pathway to trace model outputs to their training data, aligning with growing regulatory focus on transparency.  [^1] EU AI Act (final text, 2024/25): High-risk systems must have technical capabilities to “provide information relevant to explain [their] output” and enable deployers to interpret the output, plus mechanisms to collect, store and interpret logs (traceability). GPAI providers must also publish a “sufficiently detailed summary of the content used for training.” These provisions directly motivate attribution to training data. Notably, internal and external citations are complementary: internal serves as a fallback when retrieval fails or is disabled, while external helps when internal knowledge is incomplete or missing (e.g, new documents).
In this work, we ask whether LLMs can be trained to perform reliable internal citations without retrieval. Our motivation is threefold: 1) When retrieval is unavailable (e.g., due to latency or infrastructure limits), the model should still be capable of producing citations on its own. 2) When retrieval is available, internal citations can act as a safeguard and complement, helping to offset retrieval noise or missing evidence. 3) In both cases, internal citations provide an added layer of explainability by linking outputs back to training data, thereby enhancing transparency. Prior work (Khalifa et al., 2024) showed early promise that during continual pretraining, models can associate facts with document identifiers. However, their study is limited to a synthetic biography dataset with uniform, single-fact citations, leaving it unclear whether the approach generalizes to the complexity of real-world documents—longer, diverse, interdependent, and variably expressed. To address this, we introduce CitePretrainBench, a benchmark emphasizing document complexity and long-form citation. Models must index both (i) real-world corpora (Wikipedia, Common Crawl, arXiv) and (ii) novel, unseen documents, testing their ability to relearn known knowledge and acquire new knowledge with attributions. In the QA phase, we include short-form tasks requiring single citations and long-form tasks requiring synthesis across multiple documents with coherent, well-cited answers.
@@
We train our model with a two-stage framework: continual pretraining to index knowledge, and instruction tuning to elicit citation behavior. At inference, citation decoding is constrained to corpus titles for verifiability. We begin with a passive indexing baseline, where document identifiers are appended to each document (Khalifa et al., 2024). This approach allows the model to copy identifiers for memorized quotes, but our benchmark reveals key limitations that do not appear in prior synthetic set-ups (Khalifa et al., 2024): 1. Complex facts $\neq$ quotes. Many real evaluation questions require the synthesis or paraphrasing of information distributed across a document. The model rarely learns to associate such non‑verbatim facts with the correct document identifier of the original text. 2. Granularity alone isn’t enough. Inserting the identifiers closer to each fact (Khalifa et al., 2024) (e.g., per sentence or paragraph) only improves slightly. but the model still fails to ground non-verbatim content. To address this, we propose Active Indexing—a strategy that goes beyond memorization and trains the model to actively incorporate identifiers across diverse contexts. During continual pretraining, we generate synthetic data that (1) restate each fact using varied linguistic forms (e.g., definitions, comparisons, summaries), and (2) train the model to either recall knowledge given a document ID or attribute the correct ID based on the fact. This yields two complementary training objectives: 1. Source → Fact (forward): Answer questions conditioned on the given document identifier, promoting internal memory retrieval and reasoning. 2. Fact → Source (backward): Predict the document identifier for a generated answer, reinforcing attribution and source grounding.
@@
On Qwen-2.5-7B/3B, Active Indexing improves citation by up to 30.2% over passive indexing. We find (1) combining forward and backward objectives is most effective, (2) citation precision benefits more from model scale than answer correctness, and (3) Proprietary models like GPT-4.1, while substantially outperforming Qwen2.5 models in answer correctness, still fall short in generating reliable citations compared to models trained with Active Indexing. In our ablation study, we identify two key reasons why Active Indexing is effective: (1) it presents facts in greater quantity and more diverse formats, and (2) it explicitly trains the model to utilize document identifiers. Moreover, Active Indexing continues to benefit from scaling, showing no signs of saturation even when the amount of augmented data reaches 16× the original corpus. It improves both the model’s ability to memorize document identifiers and to generalize their usage to downstream tasks. Finally, we show that internal and external citations are complementary: internal excels under poor retrieval, external under strong retrieval. Our hybrid approach combines both sources to achieve the best overall performance across conditions, offering robust citations under common retrieval imperfections (Wang et al., 2025a).
@@ 2 Related Work
Attribution via Retrieval-Augmented Generation (RAG). A common approach to knowledge attribution in LLMs is to generate citations from evidence retrieved at inference time (external citations) (Rashkin et al., 2023; Yue et al., 2023; Liu et al., 2023), either before generation (Nakano et al., 2021; Menick et al., 2022b; Ye et al., 2024; Kamalloo et al., 2023) or afterward (He et al., 2023; Gao et al., 2023b). While effective, the retrieval pipeline of external citations adds computational overhead and can miss or conflict with the model’s parametric knowledge, leading to inconsistencies with what the model actually knows (Petroni et al., 2020; Xie et al., 2024; Huang et al., 2024a; Chuang et al., 2025; Huang et al., 2025). In contrast, our approach enables direct attribution to internal sources. This reduces overhead, improves explainability, and avoids dependence on noisy retrieval.
Generative Retrieval (GR) is another type of retriever used in RAG, replacing embedding-based retrieval with a generative model that maps queries directly to document IDs (Tay et al., 2022; Li et al., 2024b, a; Askari et al., 2024; Li et al., 2023). Then a separate QA model answers using the retrieved documents. Despite ID generation, retrieval and answering remain separate steps: the retriever can’t answer questions, and the QA model operates in an open-book setting, relying on external documents. As the knowledge is not internalized to QA model, this still constitutes external citation. In contrast, internal citation unifies retrieval and answering in a single closed-book model. The LLM internalizes both document IDs and their associated knowledge, enabling end-to-end answer generation with citations without external context. This is strictly harder: beyond just learning a query → docID mapping as in GR (Wang et al., 2022; Ren et al., 2023), the model must acquire knowledge and learn to use and cite it appropriately when generating answers.
Internal Knowledge and Memorization. Several studies have examined LLMs’ ability to memorize and recall training data for citation (Agrawal et al., 2024; Zuccon et al., 2023; Carlini et al., 2021), using prompting (Sun et al., 2023; Weller et al., 2024), constrained decoding (Wang et al., 2024), or fine-tuning (Zhang et al., 2025b) to improve attribution post-training. However, these approaches lack structured attribution during pretraining. Gao et al. (2025) show that prepending lightweight source hints during pretraining can steer model behavior, though it does not induce citation abilities. Source-aware training enable citations by attaching document IDs to continual-pretraining data (Khalifa et al., 2024), though has been limited to synthetic corpora, restricting methods and findings generalizability. More related work is in Appendix B.
@@ 3.1 Problem Formulation
We study internal citations: a closed-book LLM must answer a question and simultaneously produce verifiable source identifiers from its training corpus in a single pass without consulting an external retriever. Let the training corpus be $\mathcal{D}=\{(c_{i},t_{i})\}_{i=1}^{N}$, where $c_{i}\in\Sigma^{*}$ is the full text of document $i$ and $t_{i}\in\mathcal{T}$ is its unique human-readable title. Given a question $q$, the model outputs $\mathcal{R}=f\theta(q)=\langle(s_{1},C_{1}),\ldots,(s_{m},C_{m})\rangle$, where each $s_{k}$ is a factual statement answering part of $q$ and each $C_{k}\subseteq\mathcal{T}$ is a set of titles whose documents entail $s_{k}$. To guarantee validity, citation decoding space is restricted to the known title set $\mathcal{T}$. We achieve internal citations through two-stage training:
Stage 1: Continual pretraining & index learning During continual pretraining the model should (1) absorb the factual knowledge in $\mathcal{D}$, and (2) learn an internal index that maps any factual span $s\subset c_{i}$ to its title $t_{i}$. To achieve this, the baseline Passive Indexing structures pretraining examples as $(c_{i},t_{i})$ pairs, where the model predicts $t_{i}$ given $c_{i}$.
Stage 2: Citation instruction tuning. We then apply instruction tuning so the model learns to output $(s_{k},C_{k})$ pairs when answering questions, producing factual content and citations jointly.
Evaluation focuses on two aspects: (i) factual correctness, assessing the accuracy and relevance of the generated statements $s_{k}$; and (ii) citation quality, measured by the precision and recall of the predicted title sets $C_{k}$.
@@ 3.2 Datasets
To study how language models index and cite documents during continual pretraining, we propose CitePretrainBench, a benchmark designed around a continual pretraining corpus with document identifiers and downstream QA tasks that require citation from this corpus. The corpus includes documents from Wikipedia, Common Crawl, and scientific papers from arXiv, reflecting common pretraining sources. We also introduce entirely novel documents (unseen during pretraining) to test the model’s ability to learn and cite new knowledge. To ensure consistent citation targets, we perform LLM-based title deduplication across the corpus, assigning unique document identifiers. In downstream tasks, models must cite using this identifier space.
We evaluate citation performance using both long-form and short-form QA tasks, each grounded in a distinct part of the corpus: ASQA: Long-form factoid QA requiring multi-document reasoning. Sources come from the 2019/08/01 Wikipedia snapshot via the KILT knowledge base (Stelmakh et al., 2022; Petroni et al., 2021). ELI5: Open-ended long-form QA from Reddit’s “Explain Like I’m Five” forum (Fan et al., 2019), with retrieved documents from the August 2019 Common Crawl, preprocessed via CCNet (Wenzek et al., 2020). SciQAG: Short-form QA grounded in scientific papers, with document titles retrieved from arXiv (Wan et al., 2024). RepliQA: Short-form QA over fictional, synthetic documents created post-training cutoff (Monteiro et al., 2024), including gold answers and document-linked questions. For full details on datasets and corpus, see Appendix A.
@@ 3.3 Metrics: Correctness
We first evaluate the generation’s informativeness and utility—that is, its correctness with respect to the question. For ASQA, We compute Exact Match Recall (Stelmakh et al., 2022), which measures the recall of correct short answers by checking whether each reference answer appears as an exact substring in the model’s output. For ELI5, we use the Claim Recall to assess the correctness of generated answers—that is, how many gold claims are supported by the answer (Fan et al., 2019). Specifically, we compute entailment scores over three sub-claims extracted from each gold answer, providing a more accurate measure of correctness. For SciQAG, we follow prior work (Wan et al., 2024) that uses LLMs (GPT-4.1 in our case) to rate answers on a 1–5 scale across multiple dimensions, with scores normalized. We adopt the Accuracy dimension, which measures how well the answer aligns with facts from the source paper, ensuring that all claims are supported by evidence. For RepliQA, we find the recall metric from (Monteiro et al., 2024) insufficiently informative and instead adopt the relaxed variant of FreshEval (Vu et al., 2024)—a lightweight auto-rater that uses few-shot prompting with an LLM (GPT-4.1 in our case) to evaluate answer correctness.
@@ 3.4 Metrics: Citation Quality
For long-form QA tasks (e.g., ASQA and ELI5), where answers contain multiple facts and lack a single gold reference, we follow Gao et al. (2023c) and use an NLI model (TrueTeacher; Gekhman et al., 2023) to check if cited documents entail the generated claims. Citation precision is the proportion of citations that support their claims; recall is the proportion of claims that are fully supported. For short-form QA tasks (e.g., SciQAG and RepliQA), each answer corresponds to a single fact and a unique gold document. We compare model citations to the gold reference, defining precision as the fraction of citations that match, and recall as 1 if the gold citation appears, 0 otherwise. This may underestimate true accuracy, as some citations may entail the answer without matching the gold. See Appendix D for details.